# Sparse juntas on the biased hypercube

### v4

@@ 6.2 Approximation by monotone DNF
Suppose that $f\colon(\{0,1\}^{n},\mu_{p})\to\{0,1\}$ is monotone and $\epsilon$-close to degree $d$, where $p\leq 1/2$. Then $f$ is $O(\epsilon)$-close to a monotone DNF $F$ which satisfies the following properties:
@@
    For every $e$, the number of minimal high-degree inputs of $F$ of weight $e$ is $O_{d}(\epsilon/p^{e})$.
Conversely, if $F$ is a monotone DNF satisfying these properties then it is $O_{d}(\epsilon)$-close to degree $d$.
@@
Suppose that $f\colon(\{0,1\}^{n},\mu_{p})\to\{0,1\}$ is monotone and $\epsilon$-close to degree $d$. Applying the direct part of Theorem 6.1, we obtain a function $g$ satisfying the stated properties which is $O(\epsilon)$-close to $f$. Applying the converse part of Theorem 6.1, we obtain a monotone Boolean function $F$ which is $O(\epsilon)$-close to $g$, and so $O(\epsilon)$-close to $f$.
@@
Thus $F$ is a monotone width $d$ DNF, whose minterms are all $y$ of weight at most $d$ such that $g(y)=1$ and $g(z)=0$ for all $z<y$. If $y=\mathbbm{1}_{S}$ is such a minterm then $g|_{S}=y_{S}$, and so $\tilde{g}(S)=1$. Therefore the minterm support of $F$ is contained in the monomial support of $g$. Since $g$ is $(d,O(1/p))$-sparse, it follows that $F$ is also $(d,O(1/p))$-sparse, proving Item i.
@@
$\Pr[g\neq F]\leq\sum_{\mathbbm{1}_{T}\in\mathcal{H}}p^{|T|}\leq\sum_{e\leq L_{d}}p^{e}\cdot O(\epsilon/p^{e})=O(\epsilon),$
@@
$\operatorname*{\mathbb{E}}[\operatorname{dist}(g,\{0,1\})^{2}]=O(\Pr[g\notin\{0,1\}])=O(\Pr[g\neq F])=O(\epsilon).$
@@
$\operatorname*{\mathbb{E}}[(g-F)^{2}]\leq 2\operatorname*{\mathbb{E}}[(g-\operatorname{round}(g,\{0,1\})^{2}]+2\operatorname*{\mathbb{E}}[(\operatorname{round}(g,\{0,1\})-F)^{2}]\leq O(\epsilon)+2\Pr[g\neq F]=O(\epsilon).$
@@ 7 Junta approximation
Von zur Gathen and Roche [GR97] considered a similar parameter $m^{\prime}_{d}$ in which the condition $P(0)\neq P(1)$ is replaced by the condition that $P$ be non-constant. They obtained the improved bound $m^{\prime}_{d}\leq d+O(d^{0.525})$, and conjectured that in fact $m^{\prime}_{d}\leq d+O(1)$. Since $m_{d}\leq m^{\prime}_{d}$, their bound also holds for $m_{d}$.
@@
Suppose that $f\colon(\{0,1\}^{n},\mu_{p})\to\{0,1\}$ is $\epsilon$-close to degree $d$, where $p\leq 1/2$. Then for some constant $M_{d}$ depending only on $d$,
@@
    $f$ is $O_{d}(\epsilon^{1/m_{d}}+p)$-close to a constant Boolean function.
@@
    $f$ is $O_{d}(\epsilon^{1/M_{d}})$-close to a Boolean degree $d$ function.
@@
Let $A\subseteq\mathbb{R}$ be a finite set. If $f(x)\in A$ for all $x$ in the domain of $f$ then we say that $f$ is $A$-valued. A $\{0,1\}$-valued function is thus the same as a Boolean function. Everything in this paper generalizes to the $A$-valued setting, once we generalize the Kindler–Safra theorem to the $A$-valued setting (this can be done in a black-box way, as shown in the full version of [DFH19], available on arXiv (1711.09428) and ECCC (TR17-180)).
@@ 7.1 Converse part
Proof of converse part.
Let $P$ be a degree $d$ polynomial such that $P(0),\ldots,P(m_{d}-1)\in\{0,1\}$ and $P(0)\neq P(1)$. For any $p\leq 1/2$ and $\epsilon\leq 1$, define
$f_{p,\epsilon}=P(y_{1}+\cdots+y_{N}),\text{ where }N=\epsilon^{1/m_{d}}/p.$
We assume that $N$ is an integer, and in particular, $\epsilon^{1/m_{d}}\geq p$.
If fewer than $m_{d}$ of $y_{1},\ldots,y_{N}$ evaluate to $1$ then $f_{p,\epsilon}\in\{0,1\}$, and so
$\Pr[f_{p,\epsilon}\notin\{0,1\}]\leq N^{m_{d}}p^{m_{d}}=\epsilon.$
In order to upgrade this to an $L_{2}$ bound, we use Lemma 3.11. Since $y_{1}+\cdots+y_{N}$ is a $(1,1/p)$-sparse junta, Lemma 3.5 shows that $f_{p,\epsilon}$ is a $(d,O(1/p))$-sparse junta. In order to show that $f_{p,\epsilon}$ is quantized, notice that we can find coefficients $c_{e}$ such that
@@
$f_{p,\epsilon}=\sum_{e\leq d}c_{e}\sum_{\begin{subarray}{c}S\subseteq[N]\\ |S|=e\end{subarray}}y_{S},$
showing that $f_{p,\epsilon}$ is $\{c_{0},\ldots,c_{d}\}$-quantized. Therefore Lemma 3.11 shows that $f_{p,\epsilon}$ is $O(\epsilon)$-close to Boolean.
@@
$\Pr[f_{p,\epsilon}=P(0)]\geq(1-p)^{N}\geq 4^{-\epsilon^{1/m_{d}}}=\Omega(1),$
@@
$\Pr[f_{p,\epsilon}=P(1)]\geq Np(1-p)^{N-1}=\Omega(\epsilon^{1/m_{d}}).$
@@
$\operatorname*{\mathbb{E}}[(f_{p,\epsilon}-c)^{2}]\geq\frac{1}{4}\Pr[f_{p,\epsilon}\in\{0,1\}\text{ and }f_{p,\epsilon}\neq\operatorname{round}(c,\{0,1\})]=\Omega(\epsilon^{1/m_{d}}).$
Choosing a sequence of $p,\epsilon$ with $\epsilon\to 0$ and $p\leq\epsilon^{1/m_{d}}$, this shows that the constant $m_{d}$ cannot be improved in Item a. ∎
@@ 7.2 Direct part
Let $f\colon(\{0,1\}^{n},\mu_{p})\to\{0,1\}$ be $\epsilon$-close to degree $d$, where $p\leq 1/2$. Applying Theorem 5.1, we obtain a function $g\colon(\{0,1\}^{n},\mu_{p})\to\mathbb{R}$ which is $O(\epsilon)$-close to $f$ and satisfies the following properties:
@@
    The number of minimal non-Boolean inputs of $g$ of weight $e$ is $O(\epsilon/p^{e})$, for all $e$.
@@
$|\mathcal{M}_{e}|=O(1+\epsilon^{1/m_{d}}/p^{e}).$
Proof.
@@
$\Pr[\mathcal{E}]\leq\Pr[g\notin\{0,1\}]\leq\Pr[g\neq f]=O(\epsilon).$
Consequently $|\mathcal{F}|=O(\epsilon/p^{em})$. Since $|\mathcal{F}|=\Omega(|\mathcal{M}_{e}|^{m})$, we conclude that $|\mathcal{M}_{e}|=O(\epsilon^{1/m}/p^{e})=O(\epsilon^{1/m_{d}}/p^{e})$. ∎
@@
$\Pr[g\neq 0]\leq\sum_{e=1}^{d}p^{e}|\mathcal{M}_{e}|=\sum_{e=1}^{d}O(p^{e}+\epsilon^{1/m_{d}})=O(p+\epsilon^{1/m_{d}}),$
and so $\Pr[f\neq 0]\leq\Pr[f\neq g]+\Pr[g\neq 0]=O(\epsilon^{1/m_{d}}+p)$.
Since $g$ is a $(d,O(1/p))$-sparse junta, Lemma 3.10 implies that $\operatorname*{\mathbb{E}}[(g-0)^{2}]=O(\epsilon^{1/m_{d}}+p)$, and so $\operatorname*{\mathbb{E}}[(f-0)^{2}]\leq 2\operatorname*{\mathbb{E}}[(f-g)^{2}]+2\operatorname*{\mathbb{E}}[(g-0)^{2}]=O(\epsilon^{1/m_{d}}+p)$. ∎
@@
$|\operatorname{supp}_{e}(g)|=O(1+\epsilon^{1/N_{d}}/p^{e}).$
Proof.
@@
Since $\Pr[g\notin\{0,1\}]\leq\Pr[g\neq f]=O(\epsilon)$, we conclude that
$|\operatorname{supp}_{e}(g)|=O(\epsilon^{1/N_{d}}/p^{e}).\qed$
@@
Proof of Item b.
Lemma 7.5 shows that for some constant $K_{d}$ depending only on $d$, we have $|\operatorname{supp}_{e}(g)|\leq K_{d}(1+\epsilon^{1/N_{d}}/p^{e})$. Write
$g=\underbrace{\sum_{e\colon\epsilon^{1/N_{d}}\leq p^{e}}g^{=e}}_{g_{J}}+\underbrace{\sum_{e\colon\epsilon^{1/N_{d}}>p^{e}}g^{=e}}_{h},$
@@
By construction, $|\operatorname{supp}_{e}(g)|\leq 2K_{d}$ for all $e$ in the first sum, and so $g_{J}$ is an $R_{d}$-junta, where $R_{d}=2(d+1)K_{d}$. Similarly, $|\operatorname{supp}_{e}(g)|\leq 2K_{d}\epsilon^{1/N_{d}}/p^{e}$ for all $e$ in the second sum, and so
$\Pr[g\neq g_{J}]=\Pr[h\neq 0]\leq\sum_{e\colon\epsilon^{1/N_{d}}>p^{e}}p^{e}\cdot O(\epsilon^{1/N_{d}}/p^{e})=O(\epsilon^{1/N_{d}}).$
It follows that $\Pr[f\neq g_{J}]\leq\Pr[f\neq g]+\Pr[g\neq g_{J}]=O(\epsilon^{1/N_{d}})$.
If $g_{J}$ is Boolean then we are done, assuming $M_{d}\geq N_{d}$. Otherwise, since $g_{J}$ is an $R_{d}$-junta, $\Pr[g_{J}\neq f]\geq\Pr[g_{J}\notin\{0,1\}]\geq p^{R_{d}}$. Since $\Pr[g_{J}\neq f]=O(\epsilon^{1/N_{d}})$, we conclude that $p=O(\epsilon^{1/R_{d}N_{d}})$. Applying Item a, this shows that $f$ is $O(\epsilon^{1/M_{d}})$-close to the constant zero function, where $M_{d}=\max(m_{d},R_{d}N_{d})$. ∎
@@
Let $\mathcal{T}$ consist of $A\setminus\{a\}$ for all $A\in\mathcal{S}$ containing $a$, and apply the induction hypothesis. Thus either (i’) $|\mathcal{T}|\leq C_{e-1,m}$, or (ii’) there exists $J$ such that if we choose $A_{1}\setminus\{a\},\ldots,A_{m}\setminus\{a\}$ at random from $\mathcal{T}$ then with probability at least $p_{e-1,m}$, all of these sets contain $J$, and are disjoint otherwise.
If case (i’) holds then $|\mathcal{S}|\leq|\mathcal{T}|/q_{e,m}\leq C_{e-1,m}/q_{e,m}$, and so case (i) holds for $C_{e,m}=C_{e-1,m}/q_{e,m}$. If case (ii’) holds then case (ii) holds for $I=J\cup\{a\}$ and $p_{e,m}=\min(1/2,q_{e,m}^{m}p_{e-1,m})$. ∎
@@
This suggests coloring all tuples of up to $d$ elements with the value of the inner sum. Each such sum is the sum of $O(1)$ monomial coefficients of $g$, which is $B_{d}$-quantized for some finite set $B_{d}$ depending only on $d$, and therefore there are $C_{d}$ colors, for some constant $C_{d}$ depending only for $d$. The hypergraph Ramsey theorem now implies that for an appropriate value of $R$ depending only on $d$, there is a subset $A^{\prime}\subseteq\{A_{1},\ldots,A_{R}\}$ of size $m_{d}$ such that all tuples of $\ell\leq d$ elements from $A^{\prime}$ have the same color $\gamma_{\ell}(A_{1},\ldots,A_{R})$.
@@
Therefore there is a choice of $\gamma_{0},\ldots,\gamma_{\ell}$ such that if we choose $A_{1},\ldots,A_{m_{d}}\in\mathcal{M}_{e}$ at random then with probability at least $p_{e,R}/\binom{R}{m_{d}}C_{d}^{d+1}$, all tuples of $\ell\leq d$ elements from $A^{\prime}=\{A_{1},\ldots,A_{m_{d}}\}$ have color $\gamma_{\ell}$. In particular, for all $m\leq m_{d}$ such that $m\geq 1$,

### v6

@@ 6.2 Approximation by monotone DNF
Suppose that $f\colon(\{0,1\}^{n},\mu_{p})\to\{0,1\}$ is monotone and $\varepsilon$-close to degree $d$, where $p\leq 1/2$. Then $f$ is $O(\varepsilon)$-close to a monotone DNF $F$ which satisfies the following properties:
@@
    For every $e$, the number of minimal high-degree inputs of $F$ of weight $e$ is $O_{d}(\varepsilon/p^{e})$.
Conversely, if $F$ is a monotone DNF satisfying these properties then it is $O_{d}(\varepsilon)$-close to degree $d$.
@@
Suppose that $f\colon(\{0,1\}^{n},\mu_{p})\to\{0,1\}$ is monotone and $\varepsilon$-close to degree $d$. Applying the direct part of Theorem 6.1, we obtain a function $g$ satisfying the stated properties which is $O(\varepsilon)$-close to $f$. Applying the converse part of Theorem 6.1, we obtain a monotone Boolean function $F$ which is $O(\varepsilon)$-close to $g$, and so $O(\varepsilon)$-close to $f$.
@@
Thus, $F$ is a monotone width $d$ DNF, whose minterms are all $y$ of weight at most $d$ such that $g(y)=1$ and $g(z)=0$ for all $z<y$. If $y=\mathbbm{1}_{S}$ is such a minterm then $g|_{S}=y_{S}$, and so $\tilde{g}(S)=1$. Therefore the minterm support of $F$ is contained in the monomial support of $g$. Since $g$ is $(d,O(1/p))$-sparse, it follows that $F$ is also $(d,O(1/p))$-sparse, proving Item i.
@@
$\Pr[g\neq F]\leq\sum_{\mathbbm{1}_{T}\in\mathcal{H}}p^{|T|}\leq\sum_{e\leq L_{d}}p^{e}\cdot O(\varepsilon/p^{e})=O(\varepsilon),$
@@
$\operatorname*{\mathbb{E}}[\operatorname{dist}(g,\{0,1\})^{2}]=O(\Pr[g\notin\{0,1\}])=O(\Pr[g\neq F])=O(\varepsilon).$
@@
$\operatorname*{\mathbb{E}}[(g-F)^{2}]\leq 2\operatorname*{\mathbb{E}}[(g-\operatorname{round}(g,\{0,1\})^{2}]+2\operatorname*{\mathbb{E}}[(\operatorname{round}(g,\{0,1\})-F)^{2}]\leq O(\varepsilon)+2\Pr[g\neq F]=O(\varepsilon).$
@@ 7 Junta approximation
Von zur Gathen and Roche [GathenR1997] considered a similar parameter $m^{\prime}_{d}$ in which the condition $P(0)\neq P(1)$ is replaced by the condition that $P$ be non-constant. They obtained the improved bound $m^{\prime}_{d}\leq d+O(d^{0.525})$, and conjectured that in fact $m^{\prime}_{d}\leq d+O(1)$. Since $m_{d}\leq m^{\prime}_{d}$, their bound also holds for $m_{d}$.
@@
Suppose that $f\colon(\{0,1\}^{n},\mu_{p})\to\{0,1\}$ is $\varepsilon$-close to degree $d$, where $p\leq 1/2$. Then for some constant $M_{d}$ depending only on $d$,
@@
    $f$ is $O_{d}(\varepsilon^{1/m_{d}}+p)$-close to a constant Boolean function.
@@
    $f$ is $O_{d}(\varepsilon^{1/M_{d}})$-close to a Boolean degree $d$ function.
@@
Let $A\subseteq\mathbb{R}$ be a finite set. If $f(x)\in A$ for all $x$ in the domain of $f$ then we say that $f$ is $A$-valued. A $\{0,1\}$-valued function is thus the same as a Boolean function. Everything in this paper generalizes to the $A$-valued setting, once we generalize the Kindler–Safra theorem to the $A$-valued setting (this can be done in a black-box way, as shown in the full version of [DinurFH2019], available on arXiv (1711.09428) and ECCC (TR17-180)).
@@ 7.1 Converse part
Proof 7.3 (Proof of converse part).
[…]
$f_{p,\varepsilon}=P(y_{1}+\cdots+y_{N}),\text{ where }N=\varepsilon^{1/m_{d}}/p.$
We assume that $N$ is an integer, and in particular, $\varepsilon^{1/m_{d}}\geq p$.
If fewer than $m_{d}$ of $y_{1},\ldots,y_{N}$ evaluate to $1$ then $f_{p,\varepsilon}\in\{0,1\}$, and so
$\Pr[f_{p,\varepsilon}\notin\{0,1\}]\leq N^{m_{d}}p^{m_{d}}=\varepsilon.$
In order to upgrade this to an $L_{2}$ bound, we use Lemma 3.16. Since $y_{1}+\cdots+y_{N}$ is a $(1,1/p)$-sparse junta, Lemma 3.5 shows that $f_{p,\varepsilon}$ is a $(d,O(1/p))$-sparse junta. In order to show that $f_{p,\varepsilon}$ is quantized, notice that we can find coefficients $c_{e}$ such that
@@
$f_{p,\varepsilon}=\sum_{e\leq d}c_{e}\sum_{\begin{subarray}{c}S\subseteq[N]\\ |S|=e\end{subarray}}y_{S},$
showing that $f_{p,\varepsilon}$ is $\{c_{0},\ldots,c_{d}\}$-quantized. Therefore, Lemma 3.16 shows that $f_{p,\varepsilon}$ is $O(\varepsilon)$-close to Boolean.
@@
$\Pr[f_{p,\varepsilon}=P(0)]\geq(1-p)^{N}\geq 4^{-\varepsilon^{1/m_{d}}}=\Omega(1),$
@@
$\Pr[f_{p,\varepsilon}=P(1)]\geq Np(1-p)^{N-1}=\Omega(\varepsilon^{1/m_{d}}).$
@@
$\operatorname*{\mathbb{E}}[(f_{p,\varepsilon}-c)^{2}]\geq\frac{1}{4}\Pr[f_{p,\varepsilon}\in\{0,1\}\text{ and }f_{p,\varepsilon}\neq\operatorname{round}(c,\{0,1\})]=\Omega(\varepsilon^{1/m_{d}}).$
Choosing a sequence of $p,\varepsilon$ with $\varepsilon\to 0$ and $p\leq\varepsilon^{1/m_{d}}$, this shows that the constant $m_{d}$ cannot be improved in Item a.
@@ 7.2 Direct part
Let $f\colon(\{0,1\}^{n},\mu_{p})\to\{0,1\}$ be $\varepsilon$-close to degree $d$, where $p\leq 1/2$. Applying Theorem 5.1, we obtain a function $g\colon(\{0,1\}^{n},\mu_{p})\to\mathbb{R}$ which is $O(\varepsilon)$-close to $f$ and satisfies the following properties:
@@
    The number of minimal non-Boolean inputs of $g$ of weight $e$ is $O(\varepsilon/p^{e})$, for all $e$.
@@
$|\mathcal{M}_{e}|=O(1+\varepsilon^{1/m_{d}}/p^{e}).$
Proof 7.6.
@@
$\Pr[\mathcal{E}]\leq\Pr[g\notin\{0,1\}]\leq\Pr[g\neq f]=O(\varepsilon).$
Consequently $|\mathcal{F}|=O(\varepsilon/p^{em})$. Since $|\mathcal{F}|=\Omega(|\mathcal{M}_{e}|^{m})$, we conclude that $|\mathcal{M}_{e}|=O(\varepsilon^{1/m}/p^{e})=O(\varepsilon^{1/m_{d}}/p^{e})$.
@@
$\Pr[g\neq 0]\leq\sum_{e=1}^{d}p^{e}|\mathcal{M}_{e}|=\sum_{e=1}^{d}O(p^{e}+\varepsilon^{1/m_{d}})=O(p+\varepsilon^{1/m_{d}}),$
and so $\Pr[f\neq 0]\leq\Pr[f\neq g]+\Pr[g\neq 0]=O(\varepsilon^{1/m_{d}}+p)$.
Since $g$ is a $(d,O(1/p))$-sparse junta, Lemma 3.14 implies that $\operatorname*{\mathbb{E}}[(g-0)^{2}]=O(\varepsilon^{1/m_{d}}+p)$, and so $\operatorname*{\mathbb{E}}[(f-0)^{2}]\leq 2\operatorname*{\mathbb{E}}[(f-g)^{2}]+2\operatorname*{\mathbb{E}}[(g-0)^{2}]=O(\varepsilon^{1/m_{d}}+p)$.
@@
$|\operatorname{supp}_{e}(g)|=O(1+\varepsilon^{1/N_{d}}/p^{e}).$
Proof 7.9.
@@
Since $\Pr[g\notin\{0,1\}]\leq\Pr[g\neq f]=O(\varepsilon)$, we conclude that
$|\operatorname{supp}_{e}(g)|=O(\varepsilon^{1/N_{d}}/p^{e}).$
@@
Proof 7.10 (Proof of Item b).
Lemma 7.8 shows that for some constant $K_{d}$ depending only on $d$, we have $|\operatorname{supp}_{e}(g)|\leq K_{d}(1+\varepsilon^{1/N_{d}}/p^{e})$. Write
$g=\underbrace{\sum_{e\colon\varepsilon^{1/N_{d}}\leq p^{e}}g^{=e}}_{g_{J}}+\underbrace{\sum_{e\colon\varepsilon^{1/N_{d}}>p^{e}}g^{=e}}_{h},$
@@
By construction, $|\operatorname{supp}_{e}(g)|\leq 2K_{d}$ for all $e$ in the first sum, and so $g_{J}$ is an $R_{d}$-junta, where $R_{d}=2(d+1)K_{d}$. Similarly, $|\operatorname{supp}_{e}(g)|\leq 2K_{d}\varepsilon^{1/N_{d}}/p^{e}$ for all $e$ in the second sum, and so
$\Pr[g\neq g_{J}]=\Pr[h\neq 0]\leq\sum_{e\colon\varepsilon^{1/N_{d}}>p^{e}}p^{e}\cdot O(\varepsilon^{1/N_{d}}/p^{e})=O(\varepsilon^{1/N_{d}}).$
It follows that $\Pr[f\neq g_{J}]\leq\Pr[f\neq g]+\Pr[g\neq g_{J}]=O(\varepsilon^{1/N_{d}})$.
If $g_{J}$ is Boolean then we are done, assuming $M_{d}\geq N_{d}$. Otherwise, since $g_{J}$ is an $R_{d}$-junta, $\Pr[g_{J}\neq f]\geq\Pr[g_{J}\notin\{0,1\}]\geq p^{R_{d}}$. Since $\Pr[g_{J}\neq f]=O(\varepsilon^{1/N_{d}})$, we conclude that $p=O(\varepsilon^{1/R_{d}N_{d}})$. Applying Item a, this shows that $f$ is $O(\varepsilon^{1/M_{d}})$-close to the constant zero function, where $M_{d}=\max(m_{d},R_{d}N_{d})$.
@@
Let $\mathcal{T}$ consist of $A\setminus\{a\}$ for all $A\in\mathcal{S}$ containing $a$, and apply the induction hypothesis. Thus, either (i’) $|\mathcal{T}|\leq C_{e-1,m}$, or (ii’) there exists $J$ such that if we choose $A_{1}\setminus\{a\},\ldots,A_{m}\setminus\{a\}$ at random from $\mathcal{T}$ then with probability at least $p_{e-1,m}$, all of these sets contain $J$, and are disjoint otherwise.
If case (i’) holds then $|\mathcal{S}|\leq|\mathcal{T}|/q_{e,m}\leq C_{e-1,m}/q_{e,m}$, and so case (i) holds for $C_{e,m}=C_{e-1,m}/q_{e,m}$. If case (ii’) holds then case (ii) holds for $I=J\cup\{a\}$ and $p_{e,m}=\min(1/2,q_{e,m}^{m}p_{e-1,m})$.
@@
This suggests coloring all tuples of up to $d$ elements with the value of the inner sum. Each such sum is the sum of $O(1)$ monomial coefficients of $g$, which is $B_{d}$-quantized for some finite set $B_{d}$ depending only upon on $d$, and therefore there are $C_{d}$ colors, for some constant $C_{d}$ depending only for $d$. The hypergraph Ramsey theorem now implies that for an appropriate value of $R$ depending only on $d$, there is a subset $A^{\prime}\subseteq\{A_{1},\ldots,A_{R}\}$ of size $m_{d}$ such that all tuples of $\ell\leq d$ elements from $A^{\prime}$ have the same color $\gamma_{\ell}(A_{1},\ldots,A_{R})$.
@@
Therefore, there is a choice of $\gamma_{0},\ldots,\gamma_{\ell}$ such that if we choose $A_{1},\ldots,A_{m_{d}}\in\mathcal{M}_{e}$ at random then with probability at least $p_{e,R}/\binom{R}{m_{d}}C_{d}^{d+1}$, all tuples of $\ell\leq d$ elements from $A^{\prime}=\{A_{1},\ldots,A_{m_{d}}\}$ have color $\gamma_{\ell}$. In particular, for all $m\leq m_{d}$ such that $m\geq 1$,
@@
Acknowledgements
We thank the anonymous reviewers for their careful reading of the manuscript, and for significantly improving some of the proofs. The second author thanks Mathews Boban for helpful discussions.
[…]